Suppose that you have taken a long position on a put option. The strike price is $125, and the option premium / price is $10. When the option expires, the value of the underlying asset is $90. What is your pay-off and profit / loss?

please show your work
Answer #1

You have purchased put option.

premium paid = 10

Now the put option buyer will have positive payoff only if the price on expiry is lower than the strike price, otherwise payoff = zero

In our case, strike price = 125 and price on expiry = 90,

so payoff = strike price - price on expiry = 125 -90 = 35

so profit = payoff - premium paid = 35 - 10 = 25

Answer : Profit of $25 per share [Thumbs up please]